MONT 104N – Modeling the Environment Student-Led Discussions December 9 – December 13

General Information

For the final week of the semester, different assigned groups from the class will be responsible for giving oral presentations, including some questions to get a class discussion started. The assigned groups and a suggested topic for each group are given below. If you wish, your group can also choose one of the articles you reported in your journals, or something similar you might have read more recently instead.

Important Note: If you do want to present on a topic of your own choice, please notify me (either by email or by coming to office hours to discuss your ideas) about the topic you want to present no later than Monday, December 2 (the day we return from Thanksgiving break). I reserve the right to ask you to think about a different topic or to come at a topic from a different angle if two groups end up asking to do the same thing.

Let me know if you need any help tracking down sources for these or other possible topics.

The assignment will be:

• To digest a timely news story involving an environmental question,

• To present what the story says, possibly including connections with things we have talked about in the course (if they exist), and then

• To get a class discussion started about the issues involved, what people are trying to do about it, what our responses should be, and so forth.

Each group will have about half a class period (25 minutes). Aim to use about half of your group’s time presenting what your source story or article said, and the rest on the followup discussion. You may use PowerPoint or GoogleSlides (which is similar to PowerPoint) for the presentation segment if you like. If you choose that option, you probably don’t want any more than 5 slides, though, so choose carefully what you put there. As a general guideline, I would say the slides should be used to present visual information (i.e. images, graphs, charts, ... ) rather than to duplicate what you are saying. Your oral presentation will explain what’s going on, tell us what to look for, and so forth. Please don’t copy and paste whole pages from your news story or article into the slides you present.

These presentations will be going on in the class meetings of Monday, December 9, Wednesday December 11, and Friday, December 13.

I’ll be posting a review sheet and practice problems for the final exam as well. The exam will be given at the published time for MWF 9:00am classes – Friday, December 20 at 8:00am.

Schedule Please take a look at this before you leave for Thanksgiving and let me know if there are any conflicts or situations that might call for changes – I’m willing to consider rearranging things as long as everyone involved is OK with it. December 9 Group 1 – Augusta, Brooke, Caroline Suggested topic:

https://www.nytimes.com/2019/11/12/climate/energy-trends-climate-change.html

Group 2 – Christian, Jack, Kate Suggested topic:

https://www.nytimes.com/interactive/2019/11/14/climate/car-ban-air-pollution.html

December 11 Group 3 – Maggie P., Margo R., Matt V. Suggested topic:

https://www.nytimes.com/interactive/2019/10/29/climate/coastal-cities-underwater.html

Group 4 – Matthew B., Ming, Owen Suggested topic:

     https://www.nytimes.com/2019/11/13/world/europe/venice-flood.html

December 13 Group 5 – Robert, Sean, Sinead, Tom Suggested topic:

https://www.nytimes.com/2019/11/14/climate/bernie-sanders-climate-change.html